# Probabilistic Approaches to The Energy Equality in Forced Surface Quasi-Geostrophic Equations

### v1

@@ 3. Existence and Exponential Tightness
Taking the $W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))$ norm to $\theta_{\varepsilon,\delta}(t)$, it follows from the convexity of the exponential function that
$\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\|\theta_{\varepsilon,\delta}\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big{\}}\lesssim J_{1}+J_{2}+J_{3}+J_{4},$
where
$J_{1}:=$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\|\theta_{\varepsilon,\delta}(0)\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big{\}},$
$J_{2}:=$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\Big{\|}\int^{\cdot}_{0}\Lambda^{2\alpha}\theta_{\varepsilon,\delta}\Big{\|}_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big{\}},$
$J_{3}:=$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\Big{\|}\int^{\cdot}_{0}\nabla\cdot\left(\theta_{\varepsilon,\delta}R^{\perp}\theta_{\varepsilon,\delta}\right)\Big{\|}_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big{\}},$
$J_{4}:=$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon^{1/2}}\|\Lambda^{2\beta}\sqrt{Q_{\delta}}W\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big{\}}.$
Due to (B.1) and the fact that
$\|\theta_{\varepsilon,\delta}(0)\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\leqslant C(T)\|\theta_{\varepsilon,\delta}(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})},$
@@
Using the Sobolev embedding $W^{1,2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))\subset W^{\sigma,2}([0,T];H^{2\alpha-l}(\mathbb{T}^{2}))$, we deduce that
$J_{2}\lesssim$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\Big{\|}\int^{\cdot}_{0}\Lambda^{\alpha}\theta_{\varepsilon,\delta}\Big{\|}_{W^{1,2}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2}))}\Big{\}}$ (3.18)
$\lesssim$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\int^{T}_{0}\|\Lambda^{\alpha}\theta_{\varepsilon,\delta}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\Big{\}}\leqslant C(\eta,T).$
@@
$\quad|\langle\nabla\cdot\left(\theta_{\varepsilon,\delta}R^{\perp}\theta_{\varepsilon,\delta}\right),\phi\rangle|$ (3.19)
$\lesssim|\langle\theta_{\varepsilon,\delta}R^{\perp}(S_{j}\theta_{\varepsilon,\delta}),\nabla\phi\rangle|+|\langle S_{j}\theta_{\varepsilon,\delta}R^{\perp}(H_{j}\theta_{\varepsilon,\delta}),\nabla\phi\rangle|+|\langle\nabla\cdot\left(H_{j}\theta_{\varepsilon,\delta}R^{\perp}(H_{j}\theta_{\varepsilon,\delta})\right),\phi\rangle|$
$\lesssim\left[\left\|S_{j}\theta_{\varepsilon,\delta}\right\|_{H^{2\beta-\alpha}(\mathbb{T}^{2})}\|\theta_{\varepsilon,\delta}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}+\|\theta_{\varepsilon,\delta}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right]\|\phi\|_{H^{3+2\beta-\alpha+\iota}(\mathbb{T}^{2})}$
$\lesssim\|\theta_{\varepsilon,\delta}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\|\phi\|_{H^{3+2\beta-\alpha+\iota}(\mathbb{T}^{2})}.$
Consequently,
$J_{3}\lesssim\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\int^{T}_{0}\|\theta_{\varepsilon,\delta}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\Big{\}}\leqslant C(\eta,T).$ (3.20)
@@
Adding (3.17), (3.18), (3.20) and (3.23) together, we complete the proof of (3.15). Finally, (3.16) follows from exponential Chebyshev’s inequality.
@@
Let $\varepsilon,\delta$ and $\theta_{\varepsilon,\delta}$ satisfy the conditions in Lemma 3.5. Fix $\sigma\in(0,1/2)$ and $l>3+2\beta-\alpha$. Then there exists $\eta_{0}>0$ such that for every $\eta\in[0,\eta_{0}]$, under the scaling regime (2.7), we have
$\varepsilon\log\mathbb{E}\Big{(}\exp\Big{\{}\frac{\eta}{2\varepsilon}\|\theta_{\varepsilon,\delta}\|_{C^{\sigma}([0,T];H^{-l}(\mathbb{T}^{2}))}^{2}\Big{\}}\Big{)}\leqslant C(\eta,T),$
@@
$\limsup_{R\rightarrow\infty}\sup_{\varepsilon\in(0,1)}\varepsilon\log\mathbb{P}\Big{(}\|\theta_{\varepsilon,\delta}\|_{C^{\sigma}([0,T];H^{-l}(\mathbb{T}^{2}))}>R\Big{)}=-\infty.$
@@
$|\langle\int^{t}_{s}\nabla\cdot\left(\theta_{\varepsilon,\delta}R^{\perp}\theta_{\varepsilon,\delta}\right)\,\mathrm{d}r,\phi\rangle|$ $\lesssim\int^{t}_{s}|\langle\nabla\cdot\left(\theta_{\varepsilon,\delta}R^{\perp}\theta_{\varepsilon,\delta}\right),\phi\rangle|\,\mathrm{d}r$
$\lesssim\int^{t}_{s}\|\theta_{\varepsilon,\delta}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\|\phi\|_{H^{l}(\mathbb{T}^{2})}\,\mathrm{d}r$
$\lesssim(t-s)^{1/2}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,\delta}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\|\phi\|_{H^{l}(\mathbb{T}^{2})}.$
It follows that
$\quad\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\Big{\|}\int^{\cdot}_{0}\nabla\cdot\left(\theta_{\varepsilon,\delta}R^{\perp}\theta_{\varepsilon,\delta}\right)\Big{\|}_{C^{\sigma}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big{\}}$
$\lesssim\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,\delta}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\Big{\}}\leqslant C(\eta,T).$
@@
Let $\varepsilon,\delta$ and $\theta_{\varepsilon,\delta}$ satisfy the conditions in Lemma 3.5. Then under the scaling regime (2.7), there exists a sequence of compact set $\{K_{n}\}_{n\geqslant 1}\subset\mathbb{X}_{\alpha,\beta}$, such that the laws $\{\mu_{\varepsilon,\delta}\}$ of $\{\theta_{\varepsilon,\delta}\}$ satisfy
$\limsup_{n\rightarrow\infty}\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,\delta}(K_{n})=-\infty.$
@@
This follows by the same argument as in [GHW24, Corollary 4.2]. In our case, $K_{n}$ can be chosen as
@@
where $\sigma\in(0,1/2)$ and $l>3+2\beta-\alpha$. Indeed, Lemma 3.3, Lemma 3.5, and Aubin-Lions lemma imply the exponential tightness of $\{\mu_{\varepsilon,\delta}\}$ in $L^{2}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2})).$ The exponential tightness in $L_{w}^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))$ is a direct conclusion of Lemma 3.3. Finally, it follows from Lemma 3.6 that $\{\mu_{\varepsilon,\delta}\}$ is exponentially tight in $C([0,T];H^{\alpha-2\beta}_{w}(\mathbb{T}^{2}))$. ∎
@@ 4. Characterization of Rate Function
In the sequel, we will provide a variational characterization of $\mathcal{I}$, which is equivalent to the expression (4.1), (4.2) and (4.3). We now define the map $\Lambda_{0}(\cdot,\cdot):C_{b}(H^{\alpha-2\beta}_{w}(\mathbb{T}^{2}))\times H^{\alpha-2\beta}(\mathbb{T}^{2})\rightarrow\mathbb{R}$ by
@@
$+\frac{1}{2}\int_{0}^{T}\left\langle R_{2}\theta,\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}t-\frac{1}{2}\int_{s}^{t}\left\langle R_{1}\theta,\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}t-\frac{1}{2}\|\Lambda^{2\beta}\varphi\|_{L^{2}([0,T];H)}^{2}.$
@@
Let the rate functions $\mathcal{I}_{0},\mathcal{I}_{dyna},\mathcal{I}$ be defined by (4.2), (4.3) and (4.1), respectively. Then for any $\theta\in\mathbb{X}_{\alpha,\beta}$ with $\mathcal{I}(\theta)<\infty$, we have
@@
According to Proposition B.1, $\mu_{\varepsilon,\delta}^{0}=\mathcal{G}(0,\varepsilon Q_{\delta}/2)$ satisfies the large deviations on $H^{\alpha-2\beta}_{w}(\mathbb{T}^{2})$ with rate function $\mathcal{I}_{0}$ under Assumption A and the scaling regime (2.7). Thus (4.8) follows from Proposition B.1 and Bryc’s lemma (see, for example, [DZ10, Theorem 4.4.2]). Next we prove (4.9). Let $F^{T}(\cdot,\theta):C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))\rightarrow\mathbb{R}$ be the linear part of $\Lambda_{1}^{T}(\cdot,\theta):$
$F^{T}(\varphi,\theta):=$ $\langle\theta(T),\varphi(T)\rangle-\langle\theta(0),\varphi(0)\rangle-\int_{0}^{T}\langle\theta,\partial_{t}\varphi\rangle\,\mathrm{d}t+\int_{0}^{T}\langle\theta,\Lambda^{2\alpha}\varphi\rangle\,\mathrm{d}t$ (4.10)
@@
In the sequel, we prove the reverse inequality of (4.12). For the case $\sup\limits_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta)=\infty$, it is clear that
$\mathcal{I}_{dyna}(\theta)\leqslant\sup\limits_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta).$
Therefore it is sufficient to discuss the case $\sup\limits_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta)<\infty$. By the definition of $\Lambda_{1}^{T}$,
@@
Let $g=g^{\theta}:=\Lambda^{2\beta}\Psi^{\theta}\in L^{2}([0,T];H)$, (4.13) implies that $\theta$ is a weak solution of (1.1) with respect to the control $g$ in the sense of Definition 4.1. Moreover,
$\|g^{\theta}\|_{L^{2}([0,T];H)}=\|\Psi\|_{L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2}))}=\|F^{T}(\cdot,\theta)\|_{(L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2})))^{\prime}}.$
Here $(L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2})))^{\prime}$ is the dual space of the Hilbert space $L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2})).$ From the definition of $F^{T}$ and $\Lambda_{1}^{T}$, we have
$\Lambda_{1}^{T}(\varphi,\theta)=F^{T}(\varphi,\theta)-\frac{1}{2}\|\Lambda^{2\beta}\varphi\|_{L^{2}([0,T];H)}^{2}.$
Thanks to the fact that $C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))$ is dense in $L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2}))$, and the calculation of the extremum of a quadratic functional on a Hilbert space, it follows that
$\sup_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta)=$ $\sup_{\varphi\in L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2}))}\Big{(}F^{T}(\varphi,\theta)-\frac{1}{2}\|\Lambda^{2\beta}\varphi\|_{L^{2}([0,T];H)}^{2}\Big{)}$
$=$ $\frac{1}{2}\|F^{T}(\cdot,\theta)\|_{(L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2})))^{\prime}}^{2}=\frac{1}{2}\|g^{\theta}\|_{L^{2}([0,T];H)}^{2}\geqslant\mathcal{I}_{dyna}(\theta).$
This completes the proof. ∎
@@
The lower semi-continuity of the $H^{\alpha-2\beta}(\mathbb{T}^{2})$-norm with respect to the weak topology of $H^{\alpha-2\beta}(\mathbb{T}^{2})$ implies that the map $\theta\mapsto\mathcal{I}_{0}(\theta(0))$ is lower semi-continuous on $\mathbb{X}_{\alpha,\beta}$. We are left with the task of proving the lower semi-continuity of $\mathcal{I}_{dyna}$ Suppose that $\theta\in\mathbb{X}_{\alpha,\beta}$ and $\{\theta_{n}\}_{n\geqslant 1}\subset\mathbb{X}_{\alpha,\beta}$ satisfy $\theta_{n}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\beta}$ as $n\rightarrow\infty$. With the help of Proposition 4.2, it is sufficient to show that for arbitrary fixed $\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))$, the map $\Lambda_{1}^{T}(\varphi,\cdot)$ is continuous with respect to the topology of $\mathbb{X}_{\alpha,\beta}$:
@@
The definition of the topology of $\mathbb{X}_{\alpha,\beta}$ implies the convergence of $\langle\theta_{n}(T),\varphi(T)\rangle$, $\langle\theta_{n}(0),\theta(0)\rangle$ and $\int_{0}^{T}\langle\theta_{n},\partial_{t}\varphi\rangle\,\mathrm{d}t$ directly. The convergence of the dissipation term $\int_{0}^{T}\langle\theta_{n},\Lambda^{2\alpha}\varphi\rangle\,\mathrm{d}t$ follows from the fact that $\theta_{n}\rightarrow\theta$ weakly in $L^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))$.

### v3

@@ 3. Existence and Exponential Tightness
Taking the $W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))$ norm to $\theta_{\varepsilon,\delta(\varepsilon)}(t)$, it follows from the convexity of the exponential function that
$\quad\varepsilon\log\mathbb{E}\exp\Big\{\frac{\eta}{\varepsilon}\|\theta_{\varepsilon,\delta(\varepsilon)}\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big\}$
$\lesssim$ $\varepsilon\log\mathbb{E}\exp\Big\{\frac{\eta}{\varepsilon}\|\theta_{\varepsilon,\delta(\varepsilon)}(0)\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big\}+\varepsilon\log\mathbb{E}\exp\Big\{\frac{\eta}{\varepsilon}\Big\|\int^{\cdot}_{0}\Lambda^{2\alpha}\theta_{\varepsilon,\delta(\varepsilon)}\Big\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big\}$
$+\varepsilon\log\mathbb{E}\exp\Big\{\frac{\eta}{\varepsilon}\Big\|\int^{\cdot}_{0}\nabla\cdot\left(\theta_{\varepsilon,\delta(\varepsilon)}R^{\perp}\theta_{\varepsilon,\delta(\varepsilon)}\right)\Big\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big\}$
$+\varepsilon\log\mathbb{E}\exp\Big\{\frac{\eta}{\varepsilon^{1/2}}\|\Lambda^{2\beta}\sqrt{Q_{\delta(\varepsilon)}}W\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big\}=:J_{1}+J_{2}+J_{3}+J_{4}.$
Due to Proposition 2.4 and the fact that
$\|\theta_{\varepsilon,\delta(\varepsilon)}(0)\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\leqslant C(T)\|\theta_{\varepsilon,\delta(\varepsilon)}(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})},$
@@
Using the Sobolev embedding $W^{1,2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))\subset W^{\sigma,2}([0,T];H^{2\alpha-l}(\mathbb{T}^{2}))$ and Lemma 3.2, we deduce that
$J_{2}\lesssim$ $\varepsilon\log\mathbb{E}\exp\Big\{\frac{\eta}{\varepsilon}\Big\|\int^{\cdot}_{0}\Lambda^{\alpha}\theta_{\varepsilon,\delta(\varepsilon)}\Big\|_{W^{1,2}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2}))}\Big\}$ (3.12)
$\lesssim$ $\varepsilon\log\mathbb{E}\exp\Big\{\frac{\eta}{\varepsilon}\int^{T}_{0}\|\Lambda^{\alpha}\theta_{\varepsilon,\delta(\varepsilon)}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\Big\}\leqslant C(\eta,T).$
@@
$\quad|\langle\nabla\cdot\left(\theta_{\varepsilon,\delta(\varepsilon)}R^{\perp}\theta_{\varepsilon,\delta(\varepsilon)}\right),\phi\rangle|$ (3.13)
$\lesssim|\langle\theta_{\varepsilon,\delta(\varepsilon)}R^{\perp}(S_{j}\theta_{\varepsilon,\delta(\varepsilon)}),\nabla\phi\rangle|+|\langle S_{j}\theta_{\varepsilon,\delta(\varepsilon)}R^{\perp}(H_{j}\theta_{\varepsilon,\delta(\varepsilon)}),\nabla\phi\rangle|+|\langle\nabla\cdot\left(H_{j}\theta_{\varepsilon,\delta(\varepsilon)}R^{\perp}(H_{j}\theta_{\varepsilon,\delta(\varepsilon)})\right),\phi\rangle|$
$\lesssim\left[\left\|S_{j}\theta_{\varepsilon,\delta(\varepsilon)}\right\|_{H^{2\beta-\alpha}(\mathbb{T}^{2})}\|\theta_{\varepsilon,\delta(\varepsilon)}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}+\|\theta_{\varepsilon,\delta(\varepsilon)}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right]\|\phi\|_{H^{3+2\beta-\alpha+\iota}(\mathbb{T}^{2})}$
$\lesssim\|\theta_{\varepsilon,\delta(\varepsilon)}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\|\phi\|_{H^{3+2\beta-\alpha+\iota}(\mathbb{T}^{2})}.$
Consequently, according to Lemma 3.2,
$J_{3}\lesssim\varepsilon\log\mathbb{E}\exp\Big\{\frac{\eta}{\varepsilon}\int^{T}_{0}\|\theta_{\varepsilon,\delta(\varepsilon)}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\Big\}\leqslant C(\eta,T).$ (3.14)
@@
Adding (3.11), (3.12), (3.14), and (3.17) together, we complete the proof of (3.9). Finally, (3.10) follows from the exponential Chebyshev’s inequality.
@@
Let $\varepsilon,\delta(\varepsilon)$ and $\theta_{\varepsilon,\delta(\varepsilon)}$ satisfy the conditions in Lemma 3.4. Fix $\sigma\in(0,1/2)$ and $l>3+2\beta-\alpha$. Then there exists $\eta_{0}>0$ such that for every $\eta\in[0,\eta_{0}]$, under the scaling regime (1.7), we have
$\varepsilon\log\mathbb{E}\Big(\exp\Big\{\frac{\eta}{\varepsilon}\|\theta_{\varepsilon,\delta(\varepsilon)}\|_{C^{\sigma}([0,T];H^{-l}(\mathbb{T}^{2}))}^{2}\Big\}\Big)\leqslant C(\eta,T),$
@@
$\limsup_{R\rightarrow\infty}\sup_{\varepsilon\in(0,1)}\varepsilon\log\mathbb{P}\Big(\|\theta_{\varepsilon,\delta(\varepsilon)}\|_{C^{\sigma}([0,T];H^{-l}(\mathbb{T}^{2}))}>R\Big)=-\infty.$
@@
$\quad|\langle\int^{t}_{s}\nabla\cdot\left(\theta_{\varepsilon,\delta(\varepsilon)}R^{\perp}\theta_{\varepsilon,\delta(\varepsilon)}\right)\,\mathrm{d}r,\phi\rangle|\lesssim\int^{t}_{s}|\langle\nabla\cdot\left(\theta_{\varepsilon,\delta(\varepsilon)}R^{\perp}\theta_{\varepsilon,\delta(\varepsilon)}\right),\phi\rangle|\,\mathrm{d}r$
$\lesssim\int^{t}_{s}\|\theta_{\varepsilon,\delta(\varepsilon)}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\|\phi\|_{H^{l}(\mathbb{T}^{2})}\,\mathrm{d}r\lesssim(t-s)^{1/2}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,\delta(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\|\phi\|_{H^{l}(\mathbb{T}^{2})}.$
Applying Lemma 3.2, it follows that
$\quad\varepsilon\log\mathbb{E}\exp\Big\{\frac{\eta}{\varepsilon}\Big\|\int^{\cdot}_{0}\nabla\cdot\left(\theta_{\varepsilon,\delta(\varepsilon)}R^{\perp}\theta_{\varepsilon,\delta(\varepsilon)}\right)\Big\|_{C^{\sigma}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big\}$
$\lesssim\varepsilon\log\mathbb{E}\exp\Big\{\frac{\eta}{\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,\delta(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\Big\}\leqslant C(\eta,T).$
@@
Let $\varepsilon,\delta(\varepsilon)$ and $\theta_{\varepsilon,\delta(\varepsilon)}$ satisfy the conditions in Lemma 3.4. Then under the scaling regime (1.7), there exists a sequence of compact set $\{K_{n}\}_{n\geqslant 1}\subset\mathbb{X}_{\alpha,\beta}$, such that the laws $\{\mu_{\varepsilon,\delta(\varepsilon)}\}$ of $\{\theta_{\varepsilon,\delta(\varepsilon)}\}$ satisfy
$\limsup_{n\rightarrow\infty}\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,\delta(\varepsilon)}(K_{n})=-\infty.$
@@
This follows by the same argument as in [GHW24_LLNS, Corollary 4.2]. In our case, $K_{n}$ can be chosen as
@@
where $\sigma\in(0,1/2)$ and $l>3+2\beta-\alpha$. Indeed, Lemma 3.2, Lemma 3.4, and Aubin-Lions lemma imply the exponential tightness of $\{\mu_{\varepsilon,\delta(\varepsilon)}\}$ in $L^{2}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2})).$ The exponential tightness in $L_{w}^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))$ is a direct conclusion of Lemma 3.2. Finally, it follows from Lemma 3.5 that $\{\mu_{\varepsilon,\delta(\varepsilon)}\}$ is exponentially tight in $C([0,T];H^{\alpha-2\beta}_{w}(\mathbb{T}^{2}))$. ∎
@@ 4. Characterization of Rate Function
In the sequel, we will provide a variational characterization of $\mathcal{I}$, which is equivalent to the expression (4.1), (4.2), and (4.3). We now define the map $\Lambda_{0}(\cdot,\cdot):C_{b}(H^{\alpha-2\beta}_{w}(\mathbb{T}^{2}))\times H^{\alpha-2\beta}(\mathbb{T}^{2})\rightarrow\mathbb{R}$ by
@@
$+\frac{1}{2}\int_{0}^{T}\left\langle R_{2}\theta,\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}t-\frac{1}{2}\int_{0}^{T}\left\langle R_{1}\theta,\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}t-\frac{1}{2}\|\Lambda^{2\beta}\varphi\|_{L^{2}([0,T];H)}^{2}.$
@@
Let the rate functions $\mathcal{I}_{0},\mathcal{I}_{dyna},\mathcal{I}$ be defined by (4.2), (4.3), and (4.1), respectively. Then for any $\theta\in\mathbb{X}_{\alpha,\beta}$ with $\mathcal{I}(\theta)<\infty$, we have
@@
According to Proposition 2.4, $\mu_{\varepsilon,\delta(\varepsilon)}^{0}=\mathcal{G}(0,\varepsilon Q_{\delta(\varepsilon)}/2)$ satisfies the large deviations on $H^{\alpha-2\beta}_{w}(\mathbb{T}^{2})$ with rate function $\mathcal{I}_{0}$ under Assumption A and the scaling regime (1.7). Thus (4.8) follows from Proposition 2.4 and Bryc’s lemma (see, for example, [DZ10_LDP, Theorem 4.4.2]). Next we prove (4.9). Let $F^{T}(\cdot,\theta):C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))\rightarrow\mathbb{R}$ be the linear part of $\Lambda_{1}^{T}(\cdot,\theta):$
$F^{T}(\varphi,\theta)$= $\langle\theta(T),\varphi(T)\rangle-\langle\theta(0),\varphi(0)\rangle-\int_{0}^{T}\langle\theta,\partial_{t}\varphi\rangle\,\mathrm{d}t+\int_{0}^{T}\langle\theta,\Lambda^{2\alpha}\varphi\rangle\,\mathrm{d}t$ (4.10)
@@
In the sequel, we prove the reverse inequality of (4.12). It is sufficient to discuss the case that $\sup\limits_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta)<\infty$. By the definition of $\Lambda_{1}^{T}$,
@@
Let $g=g^{\theta}:=\Lambda^{2\beta}\Psi^{\theta}\in L^{2}([0,T];H)$, (4.13) implies that $\theta$ is a weak solution of (1.1) with respect to the control $g$ in the sense of Definition 4.1. Moreover, a dual Hilbert space analysis as in [GH23, Appendix B] implies that
$\sup_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta)=\frac{1}{2}\|g^{\theta}\|_{L^{2}([0,T];H)}^{2}\geqslant\mathcal{I}_{dyna}(\theta).$
@@
The lower semi-continuity of the $H^{\alpha-2\beta}(\mathbb{T}^{2})$-norm with respect to the weak topology of $H^{\alpha-2\beta}(\mathbb{T}^{2})$ implies that the map $\theta\mapsto\mathcal{I}_{0}(\theta(0))$ is lower semi-continuous on $\mathbb{X}_{\alpha,\beta}$. We are left with the task of proving the lower semi-continuity of $\mathcal{I}_{dyna}$. Suppose that $\theta\in\mathbb{X}_{\alpha,\beta}$ and $\{\theta_{n}\}_{n\geqslant 1}\subset\mathbb{X}_{\alpha,\beta}$ satisfy $\theta_{n}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\beta}$ as $n\rightarrow\infty$. With the help of Proposition 4.2, it is sufficient to show that for arbitrary fixed $\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))$, the map $\Lambda_{1}^{T}(\varphi,\cdot)$ is continuous with respect to the topology of $\mathbb{X}_{\alpha,\beta}$:
@@
The definition of the topology of $\mathbb{X}_{\alpha,\beta}$ implies the convergence of $\langle\theta_{n}(T),\varphi(T)\rangle$, $\langle\theta_{n}(0),\theta(0)\rangle$, and $\int_{0}^{T}\langle\theta_{n},\partial_{t}\varphi\rangle\,\mathrm{d}t$ directly. The convergence of the dissipation term $\int_{0}^{T}\langle\theta_{n},\Lambda^{2\alpha}\varphi\rangle\,\mathrm{d}t$ follows from the fact that $\theta_{n}\rightarrow\theta$ weakly in $L^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))$.